OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x^2 – 6x – 7. Find f(g(2))

OpenStudy (gavlois1):

The notation f(g(2)) means that for every x in f(x), we replace it with g(2). This means we need to compute g(2) first, which is achieved by substituting 2 for every x in g(x). \[g(x) = x^2 - 6x -7\] \[g(2) = 2^2 - 6(2) -7\] \[g(2) = 4-12-7\] From this we see that g(2) is equal to -15, and we replace every x in f(x) with -15. \[f(x) = x+8\] \[f(g(2)) = f(-15) = -15 +8\] Therefore, f(g(2)) is equal to -7

OpenStudy (anonymous):

thank you :)

OpenStudy (gavlois1):

You're welcome. Good luck with the rest of of your assignment.